OpenStudy (callisto):

sorry for asking again, when should i use disc method and shell method when i need to find the volume by integration?

OpenStudy (amistre64):

whenever you feel that one is easier than the other to perform

OpenStudy (callisto):

let say, the question is: Find the volume of the solid of revolution generated by revolving the region bounded by the curve y=xlnx, the x axis and the line x=1 and x=3 about the y-axis. what should i do?

OpenStudy (amistre64):

draw a picture :)

OpenStudy (amistre64):

\[\int_{1}^{3}2pi(xlnx)dx\] might be simplest

OpenStudy (amistre64):

forgot an x in that ...

OpenStudy (amistre64):

2pix * f(x)

OpenStudy (amistre64):

or if es are better for you; convert it to es and play along

OpenStudy (callisto):

so you use shell method, but why?

OpenStudy (amistre64):

becasue it works?

OpenStudy (amistre64):

what would y=x lnx look like if you solve for x

OpenStudy (amistre64):

y = ln x^x e^y = x^x realy isnt a simple function to play with

OpenStudy (amistre64):

so the shel method would be your simplest route to take

OpenStudy (callisto):

so the curve y=xlnx revolved about y axis = use shell method?

OpenStudy (amistre64):

it equals: use the easiest method. That method just so happens to be the shell, yes

OpenStudy (callisto):

a disc and a shell is really similar..

OpenStudy (amistre64):

they are similar, except for the way the process looks at the function

OpenStudy (amistre64):

the disc in this case looks at x(y) the shell in this case looks at y(x)

OpenStudy (amistre64):

x(y) is pretyy hard to define when y(x) = xlnx

OpenStudy (callisto):

x(y) =?

OpenStudy (amistre64):

dunno, since e^y = x^x we cant really get an explicit form for x(y)

OpenStudy (callisto):

What does x(y) mean?

OpenStudy (amistre64):

the same thing any function notation means; x as a function of y

OpenStudy (amistre64):

y(x) means y as a function of x r(t) means r as a function of t etc .....

OpenStudy (amistre64):

x(y) in this case would need to be the inverse of y(x)

OpenStudy (callisto):

so if it can be easily expressed in terms of y, disc method can also be used?

OpenStudy (amistre64):

yes, if it could be easily expresed in terms of y, then disc would be fine to use

OpenStudy (amistre64):

the method doesnt change the value

OpenStudy (amistre64):

which is more; 5 gallons of water in a blue vase; or 5 gallons of water in a purple vase?

OpenStudy (callisto):

blue vase? purple case?

OpenStudy (callisto):

express the equation into an easier form to integrate when y is expressed in terms of x, and it is revolved about y axis, shell method is used when x is expressed in terms of y, and it is revolved about y axis, disc method is used can i conclude in this way?

OpenStudy (amistre64):

in general, no

OpenStudy (amistre64):

the disc method is blocked by this problem simply becasue there is no way to define x(y) in a usable maner to play with the disc method

OpenStudy (amistre64):

the only option we have is to use the shell method

OpenStudy (callisto):

then how would you determine which is the eaiser way?

OpenStudy (callisto):

easier

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=inverse+y%3Dx*lnx that W function they use aint something that can be integrated with the stuff we are taught

OpenStudy (callisto):

i think i understand it a little, thanks :)